OpenStudy (anonymous):

please help!!!!! Find cos O If sin 0 = 0.56.

OpenStudy (anonymous):

\[\sin^2 \theta + \cos^2 \theta = 1\]

OpenStudy (anonymous):

\[\cos \theta = \sqrt{1 - \sin^2 \theta}\]

OpenStudy (anonymous):

\[\cos \left(\sin ^{-1}(0.56)\right)=0.828493 \]

OpenStudy (anonymous):

That's what I got too

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

I guess robotobey's solution is better :P